### v1

# MathEDU: Towards Adaptive Feedback for Student Mathematical Problem-Solving

Abstract
Online learning enhances educational accessibility, offering students the flexibility to learn anytime, anywhere. However, a key limitation is the lack of immediate, personalized feedback, particularly in helping students correct errors in math problem-solving. Several studies have investigated the applications of large language models (LLMs) in educational contexts. In this paper, we explore the capabilities of LLMs to assess students’ math problem-solving processes and provide adaptive feedback. The MathEDU dataset is introduced, comprising authentic student solutions annotated with teacher feedback. We evaluate the model’s ability to support personalized learning in two scenarios: one where the model has access to students’ prior answer histories, and another simulating a cold-start context. Experimental results show that the fine-tuned model performs well in identifying correctness. However, the model still faces challenges in generating detailed feedback for pedagogical purposes.
MathEDU: Towards Adaptive Feedback for Student Mathematical Problem-Solving
Wei-Ling Hsu, Yu-Chien Tang, An-Zi Yen Department of Computer Science, National Yang Ming Chiao Tung University, Taiwan weiling.hsu.cs11@nycu.edu.tw, tommytyc.cs10@nycu.edu.tw, azyen@nycu.edu.tw
1 Introduction
In the post-pandemic era, online learning has emerged as one of the mainstream methods of education (Alqahtani and Rajkhan, 2020; Jafar et al., 2022). Online educational platforms enable students to study anytime, anywhere, offering extensive question banks to assess understanding. However, without immediate teacher support, students may struggle to correct mistakes if they lack a clear grasp of underlying concepts.
Recent advancements in LLMs have prompted numerous studies (Tack and Piech, 2022; Kasneci et al., 2023) to explore their applications in the field of education. Some works (Shen et al., 2021; Yu et al., 2021; Jie et al., 2022) have explored the math word problem-solving capabilities of LLMs, yielding promising results. Leveraging the advanced natural language understanding and generation capabilities of LLMs, we explore their potential in mathematical problem-solving by focusing on the generation of free-text rationalizations.
For effective assessment of student performance, the LLM must possess a robust understanding of mathematical concepts to identify and analyze errors accurately. Additionally, it should provide clear explanations and deliver constructive, adaptive feedback tailored to the student’s reasoning. Yen and Hsu (2023) have employed GPT-3.5 (Ouyang et al., 2022) to simulate both student and teacher roles in order to examine the model’s behavior and grading capabilities. While LLMs can generate comprehensive explanations, findings indicate that they often struggle to accurately interpret students’ mathematical problem-solving processes, leading to incorrect assessments.
To better understand these challenges and explore ways to improve LLMs’ evaluative capabilities, we collect real-world data on student problem-solving processes alongside teacher-written feedback. We extend the existing MathQA (Amini et al., 2019), which contains GRE-level questions requiring advanced mathematical knowledge, by incorporating detailed annotations of students’ problem-solving processes and corresponding teacher feedback. We invited six students to solve these problems, and their answers were reviewed and graded by three mathematics experts. This process resulted in a comprehensive dataset of 4,048 annotated entries. Table 1 provides an example of the annotations, including error types (e.g., incorrect operations), incorrect steps, and teacher feedback addressing misunderstandings. Details of the dataset construction are provided in the following section.
Problem: Two trains of equal length are running on parallel lines in the same directions at 46 km/hr and 36 km/hr. The faster train passes the slower train in 144 seconds. The length of each train is:
Student Process: $46-36=10$ $10\,\mathrm{km/hr}\times\frac{5}{18}=\frac{100}{36}\,\mathrm{m/s}$ $\frac{100}{36}\times 144=400\,\mathrm{m}$
Error Type: Wrong Mathematical Operation/Concept
Error Equation: $\frac{100}{36}\times 144=400\,\mathrm{m}$
Teacher Feedback: To overtake the other train, you need to travel the combined length of both trains. Since both trains are of the same length, you need to divide by 2 to get the answer.
Table 1: Example of Student Problem-Solving Process with Teacher Grading and Feedback.
Rather than focusing solely on whether the final answer matches a predefined solution, this paper investigates how LLMs can evaluate students’ reasoning, recognizing multiple valid approaches. To facilitate this exploration, the task was divided into three subtasks: (1) answer accuracy assessment, (2) error identification, and (3) feedback generation. We investigate the capabilities of LLMs using both direct prompting and fine-tuning approaches. During fine-tuning, we employed three training strategies: single-task training, multi-task training, and end-to-end training. To evaluate the model’s performance in real-world scenarios, experiments were conducted under two conditions: one with access to a student’s prior answer history and another simulating a cold-start scenario without any prior data. In sum, our contributions are threefold: (1) This study investigates the challenge of assessing students’ reasoning processes by exploring the use of LLMs in the context of mathematical education. (2) We present the MathEDU dataset, [^1] The dataset will be released upon acceptance. An anonymized version is available at: https://anonymous.4open.science/r/MathEDU-4628/ designed to address students’ mistakes in math word problem-solving and provide personalized adaptive feedback. (3) We evaluate LLM performance with prompting and LoRA Hu et al. (2021) fine-tuning. Experimental results show that fine-tuning improves accuracy in identifying correctness but remains limited in generating adaptive feedback.
2 Related Work
2.1 Large Language Models for Education
Teachers play a crucial role in identifying and explaining student errors, significantly enhancing understanding and learning efficiency (Robinson and Loeb, 2021; Zhang et al., 2023). To address the lack of interaction and feedback in online learning, researchers have explored automated systems, such as integrating Cognitive Task Analysis (CTA) with LLMs to support student remediation (Graesser et al., 2004; Wang et al., 2024). However, the resource-intensive nature of CTA, requiring expert involvement and complex step design, limits its scalability in routine educational contexts. Some studies leverage LLMs to support instruction by generating distractors for multiple-choice questions (Dave et al., 2021), analyzing students’ answer histories (Gao et al., 2021), and creating personalized instructional materials (He-Yueya et al., 2024). Others use LLMs to simulate students in interactive chats, enhancing classroom simulations for teacher training (Markel et al., 2023). LLMs have also been deployed as virtual tutors, such as generating code explanations in computer science classrooms (MacNeil et al., 2023).
2.2 Mathematics Education Datasets
Table 2 presents a comparison of existing datasets. The ASSISTments datasets, [^2] https://www.etrialstestbed.org/resources/featured-studies/dataset-papers collected from the free online tutoring platform ASSISTments, focus on addressing the knowledge tracing problem. They consist of grade school math exercises, featuring various question types, including multiple choice, text, and open-ended questions. Multiple versions of the datasets have been released, with data collected during different periods. The Algebra 2005-2006 and Algebra 2006-2007 datasets, collected from the Cognitive Tutors system, capture student responses to Algebra problems.
Dataset #Students #Questions #Answers Solving Process Teacher’s Feedback
ASSISTments 2009 4,217 26,688 346,860
ASSISTments 2009 46,674 179,999 6,123,270
ASSISTments 2015 19,917 - 708,631
ASSISTments 2017 1,709 3,162 942,816
Algebra2005-2006 575 1,084 813,661
Algebra2006-2007 1,840 90,831 2,289,726
Bridge to Algebra 1,146 19,258 3,686,871
EdNet-KT1 784,309 13,169 95,293,926
EdNet-KT2 297,444 13,169 56,360,602
EdNet-KT3 297,915 13,169 89,270,654
EdNet-KT4 297,915 13,169 131,441,538
Junyi 72,630 25,785 16,217,311
Eedi 118,971 27,613 15,867,850
MATHDIAL - 2,861 2,861 $\checkmark$ $\checkmark$
MathEDU 6 4,048 4,048 $\checkmark$ $\checkmark$
Table 2: Comparison of Existing Datasets.
The EdNet dataset (Choi et al., 2020) is a hierarchical dataset designed for educational research. It consists of four levels (KT1 to KT4) that progressively capture more detailed student interactions. These range from question-solving logs to comprehensive action sequences, including behaviors like watching lectures or reading materials. The Junyi dataset (Pojen et al., 2020) consists of student exercise attempt logs collected during 2018-2019. It focuses on high school level Math exercise practice in Taiwan. The Eedi dataset (Wang et al., 2020), collected students’ responses to questions between 2018 and 2020, spans from primary to high school levels. It supports tasks such as predicting student answer correctness and includes additional tasks for assessing question quality and generating personalized question sequences.
Previous datasets for knowledge tracing typically span over a year of diverse student behavioral data, providing valuable insights into student engagement. However, they lack detailed records of students’ step-by-step problem-solving processes in mathematics. Furthermore, they do not include teacher feedback on student errors. To address the limitation, the MATHDIAL dataset (Macina et al., 2023) utilizes LLMs to simulate student responses in mathematical problem-solving, with teachers providing targeted instructional feedback to create educational dialogues.
Although LLMs are capable of simulating student responses, some studies (Aher et al., 2023; Markel et al., 2023) have indicated that these simulated answers are often unrealistic, particularly in maintaining a consistent knowledge level. Consequently, collecting real student problem-solving processes along with the corresponding teacher feedback is essential.
All General Gain Physics Geometry Probability Other Avg. Equations/Problem Avg. Words/Problem
Student 1 70.57% 67.75% 72.73% 74.73% 69.39% 28.57% 72.97% 3.9516 0.6442
Student 2 87.59% 86.53% 87.22% 88.83% 87.18% 100.00% 91.67% 7.3124 6.7532
Student 3 71.09% 70.11% 65.44% 77.78% 64.29% 42.86% 81.25% 5.5958 5.0943
Student 4 75.30% 77.24% 66.43% 77.84% 78.26% 87.50% 77.42% 6.7212 4.0303
Student 5 67.01% 66.55% 61.15% 73.12% 65.85% 75.00% 60.61% 4.9237 2.8871
Student 6 80.61% 77.48% 78.74% 88.00% 72.73% 85.71% 85.29% 4.1484 0.3106
Table 3: Correctness and Problem-Solving Characteristics of Students.
3 Dataset construction
3.1 Collection of Real-World Student Math Problem-Solving Processes
We construct the MathEDU dataset based on MathQA. [^3] The dataset selection criteria are detailed in Appendix A. We removed the multiple-choice options provided by MathQA, presenting the questions as open-ended tasks. To comprehensively assess students’ mathematical abilities and collect problem-solving processes from individuals with varying skill levels, we recruited six students from different university departments. Before the annotation process, students were asked to attempt 10 questions to ensure their responses adhered to the required format. This step helped avoid issues such as incomplete equations or disorganized layouts, facilitating smoother data collection.
To ensure that each student answered a balanced number of questions across all types (i.e., general, gain, physics, geometry, probability, and other), we sampled 4,500 unique questions from the six types in MathQA, maintaining the proportional distribution of the original dataset. Each student was assigned 750 questions, with the distribution of question types consistent across students. We found that the dataset contained errors such as unrecognizable symbols, unclear problem definitions, and missing accompanying images, which could hinder accurate interpretation. Therefore, students were allowed to note these issues and skip such questions. The missing questions were then supplemented until each student completed 750 questions with problem-solving processes.
Students wrote their problem-solving processes on paper, and their responses were then transcribed into format [^4] https://mathpix.com/ to ensure machine readability. We requested problem-solving steps to be as complete as possible to ensure that the teachers grading their answers could clearly understand their reasoning. If needed, they could also provide verbal explanations to elaborate on their reasoning. For questions they could not solve, students were required to note the reasons. If their handwriting was unclear, we asked them to revise to ensure legibility for scanning and conversion to format. [^5] The annotation guidelines are elaborated in Appendix B.1.  [^6] The question distribution and academic backgrounds of student annotators are detailed in the Appendix B.2.
3.2 Annotation for Error Identification and Teacher Feedback
After collecting the students’ problem-solving results, we proceeded to label detailed grading information for each answer. We invited three experts in mathematics education to serve as teachers for grading and reviewing student responses. The three invited mathematics experts annotated the specific steps where errors occurred and providing explanations to clarify the reasons for these mistakes. They also categorized the errors based on these annotated steps. Inspired by Wijaya et al. (2014), We define three categories. The error categories include “Wrong Mathematical Operation/Concept”, “Calculation Error” and “Incomplete Answer”. Additionally, to address instances where students either made careless mistakes or were completely unable to answer a question, we introduced two additional error categories: “Careless Error” and “Lack of Necessary Mathematical Concepts”. [^7] Descriptions and examples of these error categories are provided in Appendix C.
Each student’s answer was annotated by three experts. In addition to labeling the types of errors and incorrect problem-solving steps, experts provided suggestions for students, including explanations for students’ mistakes and suggestions for correcting them. The questions are reviewed and removed erroneous ones that led to unrecognizable problem-solving processes, resulting in a dataset of 4,048 real student answers. To assess the quality of the annotations, we calculated the inter-rater agreement for error category annotations, using Krippendorff’s Alpha. The result was 0.7818, indicating a substantial level of agreement among the experts in their annotations. When a disagreement occurs, we follow the majority rule to determine the final annotation. In cases where a majority decision could not be reached, the instance was revisited by the three experts to reach a final consensus. Finally, authors with expertise in mathematics education reviewed and refined the labeled data to ensure quality. As a result, our dataset includes 3,050 correct answers and 998 incorrect ones.
4 Data Analysis
4.1 Student Mathematical Abilities
Table 3 presents the correctness of students’ problem-solving processes, highlighting the diverse data we collected even from just six participants. Correctness varied significantly, with the most proficient students achieving up to 87.59%, while those with lower proficiency scored as low as 67.01%. Furthermore, individual strengths differed across areas, with some students reaching 100% accuracy in probability questions, whereas others struggled, achieving only 28.57%.
4.2 Student Problem-Solving Styles
We further analyze variations in students’ problem-solving strategies. Since quantifying these strategies is challenging, we indirectly assess them by evaluating the number of mathematical expressions used by each student. Table 3 shows the average use of equations and written explanations per problem, presented in the last two columns, providing insights into their approaches. This analysis reveals that Students 1, 5, and 6 tend to have more concise problem-solving processes, while Students 2, 3, and 4 have more detailed and complex processes. Notably, Student 2 used an average of 7.3124 equations per problem, the highest among the group.
We also analyzed the average number of English words used, excluding algebraic symbols and expressions, as students could provide verbal explanations to support their solutions. The results reveal significant differences in text usage: Student 6 rarely used words, averaging 0.3106 per problem, while Student 2 provided detailed explanations, averaging 6.7532 words per problem.
5 Methodology
In this section, we focus on how the tutor model is developed to analyze students’ problem-solving processes. The model identifies specific steps where errors occur in incorrect answers and provides targeted feedback. The $i$-th input data consists of the math word question $q_{i}$, the rationale $r_{i}$ provided in MathQA, and the student’s answering process $s_{i}$, which includes both equations and written descriptions of their reasoning. Based on this input, the model performs three main tasks:
Answer Accuracy Assessment: This task determines whether the student’s overall answer is correct based on their problem-solving process, treated as a binary classification with the output label $c_{i}$.
Problem-Solving Error Identification: This task identifies incorrect equations in the process. The output $E_{i}={e_{i,1},e_{i,2},\ldots,e_{i,\alpha}}$ consists of $\alpha$ identified incorrect equations.
Feedback Generation: This task generates feedback, including explanations or suggestions, to help students understand and correct their mistakes. The feedback is denoted as $T_{i}$.
We employ LoRA to fine-tune an LLM. [^8] The details of methods are reported in Appendix D. Three fine-tuning settings are explored. In single-task training, separate models were fine-tuned for each of the three subtasks. In multi-task training, a single model was trained on data annotated for all three subtasks, using distinct prompts to guide the model in performing specific tasks. End-to-end training required the model to sequentially analyze the problem statement and evaluate the student’s answer, providing a complete assessment without dividing the tasks explicitly.
6 Experiments
6.1 Experimental Setup
We examined the performances of Llama3 8B (Touvron et al., 2023), Llama3 70B, and GPT-3.5 with few-shot prompting. [^9] The version we used is gpt-3.5-turbo-0125. We further fine-tuned Llama3 8B with a learning rate set to 2e-4 and a LoRA rank of 16.101010All the prompts are shown in Appendix E. During inference, the temperature was set to 0, with a maximum output of 512 tokens, while other parameters remained at their default settings.
Data Splitting. To evaluate the model’s performance across different application scenarios, we designed two dataset splitting settings. (1) The first setting employs a time-series split, simulating a situation where the model has access to a student’s prior answer history. Student records were organized chronologically, with part of the records serving as the training set (answer history) and the remaining records as the validation and test sets in a 70:15:15 ratio. This approach assesses whether the model can leverage past responses to identify a student’s strengths and weaknesses and apply this knowledge to grade new answers. (2) The second setting also uses a time-series split but applies a leave-one-out strategy to simulate the model encountering a new student without prior history. In this setup, one student’s data is designated as the test set, while the data from the remaining students is proportionally divided into training and validation sets in a 9:1 ratio. This scenario, akin to the cold-start setting, evaluates the model’s adaptability to diverse problem-solving styles. [^11] The data splitting statistics are in Appendix F.
Inference Prompt. For the few-shot prompting, six examples were provided, consisting of three correct and three incorrect student responses. In the time-series split setting, we randomly selected the student’s prior answer records from the training set as few-shot demonstrations. In the leave-one-out setting, few-shot demonstrations were randomly sampled from the answer records of other students. For the fine-tuned model, which was extensively trained on the task, we employed zero-shot prompting to generate results directly.
Time Series Split Leave-One-Out
All General Gain Physics Geo. Prob. Other All General Gain Physics Geo. Prob. Other
Llama3 8B w/o $r$ 62.02% 64.66% 64.86% 55.56% 77.50% 62.50% 48.48% 63.53% 66.57% 64.61% 58.50% 69.21% 55.06% 50.31%
w/ $r$ 70.15% 71.55% 68.47% 68.42% 80.00% 75.00% 60.61% 68.58% 71.59% 68.02% 63.81% 75.47% 61.31% 61.45%
Llama3 70B w/o $r$ 83.58% 82.76% 88.29% 79.53% 97.50% 56.25% 90.91% 85.25% 86.54% 87.90% 80.89% 87.53% 74.70% 84.92%
w/ $r$ 81.76% 78.88% 87.39% 79.53% 97.50% 62.50% 84.85% 84.12% 84.57% 87.92% 81.01% 82.66% 63.69% 85.29%
GPT-3.5 w/o $r$ 40.96% 40.95% 45.05% 38.01% 52.50% 50.00% 24.24% 41.70% 43.45% 46.03% 36.91% 41.67% 37.50% 32.76%
w/ $r$ 50.41% 50.00% 54.95% 49.12% 57.50% 43.75% 39.39% 49.10% 51.65% 52.90% 43.61% 48.25% 59.23% 37.68%
Single-task w/o $r$ 87.23% 85.77% 90.09% 88.30% 90.00% 68.75% 87.88% 71.93% 71.43% 72.00% 71.98% 72.51% 72.32% 73.62%
w/ $r$ 92.70% 90.09% 95.50% 92.98% 95.00% 87.50% 100.00% 92.45% 92.45% 93.32% 91.73% 92.42% 88.99% 93.57%
Multi-task w/o $r$ 73.13% 70.26% 71.17% 80.70% 67.50% 56.25% 75.76% 43.07% 42.41% 41.95% 45.03% 44.03% 40.48% 41.91%
w/ $r$ 78.11% 74.57% 80.18% 83.63% 75.00% 62.50% 78.79% 54.77% 53.71% 53.87% 55.78% 58.00% 64.88% 55.99%
End-to-end w/o $r$ 87.73% 84.05% 93.69% 90.06% 90.00% 62.50% 90.91% 82.91% 82.29% 83.80% 83.67% 83.42% 75.00% 80.15%
w/ $r$ 91.54% 88.79% 95.50% 93.57% 95.00% 62.50% 96.97% 93.13% 93.01% 93.31% 92.53% 93.97% 95.83% 94.61%
Table 4: Results of Answer Accuracy Assessment.
6.2 Experimental Results
Answer Accuracy Assessment. We evaluated the ability of LLMs to determine the correctness of student answers. Accuracy is adopted as the evaluation metric. The results are shown in Table 4. We tested whether the detailed rationales in MathQA assist LLMs in grading student answers. “w/o $r$” and “w/ $r$” refer to whether the input includes the rationale of each question, respectively. In the time-series split setting, Llama3 70B achieved 83.58% accuracy, significantly outperforming Llama3 8B (62.02%) by 21.56%, highlighting the importance of parameter size in mathematical computation and understanding. GPT-3.5’s lower accuracy stemmed from its strict evaluation of omitted calculations. For example, if a student’s final answer was correct but some intermediate steps were skipped, GPT-3.5 often classified the answer as incorrect.
The bottom three rows of Table 4 represent our fine-tuned models, which demonstrate higher accuracy in determining correctness compared to the original, non-fine-tuned Llama3 8B model. Among these, the end-to-end training method perform better than the multi-task training method. This improvement may be attributed to the end-to-end approach, which requires the model to generate grading results sequentially, enabling it to perform more coherent reasoning and achieve better outcomes.
The impact of including rationales in the input is evident in the results shown in Table 4. For instance, Llama3 8B’s accuracy increased from 62.02% to 70.15%. Notably, Llama3 70B exhibited a slight decrease in accuracy from 83.58% to 81.76%. In general, the inclusion of rationales allowed most models to better assess correctness by providing additional contextual support, though the ability to interpret and utilize these rationales varied across models. Fine-tuned models, particularly the end-to-end model, demonstrated a stronger ability to effectively leverage rationales
In the leave-one-out setting, non-fine-tuned LLMs achieved accuracy consistent with the time-series split. In contrast, fine-tuned multi-task models saw a 20% accuracy drop and struggled with unseen problem-solving approaches, occasionally failing to produce outputs. However, the end-to-end model remained stable, achieving the highest accuracy of 93.13%, adapting effectively to solving styles excluded from training. We further conducted McNemar’s test to assess the significance of differences between our fine-tuned models and the baseline Llama3 70B without rationale. In the time-series split setting, the single-task model with rationale significantly outperformed the baseline, with $p<0.001$. Similarly, in the leave-one-out setting, the end-to-end model with rationale also achieved significantly better performance than the baseline, with $p<0.001$.
Although the end-to-end model seems to perform worse in the time-series split setting than in the leave-one-out setting, this likely reflects differences in data distributions, making direct comparison invalid. Moreover, the lack of user embeddings and effective encoding of individual records may limit the model’s ability to capture answering patterns, which we leave as future work.
Time Series Split Leave-One-Out
EM ($\uparrow$) Distance ($\downarrow$) EM ($\uparrow$) Distance ($\downarrow$)
Llama3 8B w/o $r$ 28.40% 68.44 10.58% 96.56
w/ $r$ 11.73% 97.73 8.80% 102.12
Llama3 70B w/o $r$ 32.10% 67.35 23.30% 78.33
w/ $r$ 29.63% 73.38 24.34% 78.55
GPT-3.5 w/o $r$ 30.25% 57.93 27.81% 63.07
w/ $r$ 27.78% 62.32 29.46% 63.26
Single-task w/o $r$ 24.69% 95.64 23.41% 96.85
w/ $r$ 23.46% 97.21 21.98% 99.09
Multi-task w/o $r$ 23.46% 95.70 22.59% 93.77
w/ $r$ 17.28% 96.35 22.11% 97.92
End-to-end w/o $r$ 30.25% 71.48 29.39% 74.01
w/ $r$ 40.12% 51.78 42.81% 52.21
Table 5: Results of Problem-Solving Error Identification.
Time Series Split Leave-One-Out
ROUGE-1 ROUGE-2 ROUGE-L BERTScore LLM Rating ROUGE-1 ROUGE-2 ROUGE-L BERTScore LLM Rating
Llama3 8B w/o $r$ 0.2593 0.0809 0.2391 0.7654 2.68 0.2103 0.0615 0.1924 0.6851 2.38
w/ $r$ 0.1819 0.0557 0.1632 0.5515 2.14 0.1738 0.0515 0.1594 0.5600 1.97
Llama3 70B w/o $r$ 0.2587 0.0881 0.2439 0.7147 3.61 0.2412 0.0761 0.2254 0.7219 3.77
w/ $r$ 0.2456 0.0876 0.2311 0.7108 3.31 0.2367 0.0720 0.2201 0.7205 3.56
GPT-3.5 w/o $r$ 0.2897 0.1041 0.2695 0.8395 3.70 0.2824 0.0978 0.2601 0.8477 3.39
w/ $r$ 0.2848 0.1025 0.2597 0.8366 3.63 0.2759 0.0925 0.2533 0.8451 3.31
Single-task w/o $r$ 0.1661 0.0730 0.1594 0.4836 0.63 0.0886 0.0359 0.0841 0.2639 0.47
w/ $r$ 0.0790 0.0378 0.0750 0.2182 0.39 0.0594 0.0205 0.0561 0.1828 0.33
Multi-task w/o $r$ 0.0130 0.0049 0.0124 0.0437 0.08 0.0600 0.0237 0.0577 0.1832 0.31
w/ $r$ 0.1650 0.0719 0.1552 0.4372 0.55 0.0244 0.0068 0.0230 0.0807 0.22
End-to-end w/o $r$ 0.1862 0.0751 0.1761 0.5327 0.99 0.1808 0.0637 0.1720 0.5726 2.14
w/ $r$ 0.2842 0.1418 0.2703 0.6786 1.91 0.2779 0.1089 0.2614 0.7940 2.28
Table 6: Results of Feedback Generation.
Problem-Solving Error Identification. To evaluate the model’s performance in identifying erroneous equations in cases where a student’s answer is incorrect, we employ two metrics. The first is the exact match (EM) ratio, which measures the correspondence between model-detected step and manually annotated step. The second metric is the distance, defined as the textual distance between the model-detected errors and the manually labeled errors. Textual distance measures the character difference between the predicted and annotated errors, reflecting the model’s prediction deviation. [^12] The details of the distance metric is in Appendix G.
The results are shown in Table 5. In the time series split setting, non-fine-tuned LLMs achieved an EM ratio of around 30% without rationale. However, providing rationale reduced the EM ratio for all models, with Llama3 8B showing the largest drop of 16.67%. This may result from differing formats between the rationale and the student’s actual process (e.g., -formatted), which increases the complexity of error identification.
For fine-tuned models, the single-task model performed the worst without rationale. It rarely identified errors correctly, often gave irrelevant responses, and sometimes failed to output answers. Its performance further declined when rationale was included. The multi-task model showed a similar pattern. In contrast, the end-to-end model outperformed Llama3 8B without rationale and showed noticeable improvements with rationale. This result indicates that fine-tuning with this approach helps the model better understand and utilize rationale for error identification.
In the leave-one-out setting, Llama3 8B with few-shot prompting performed poorly. Other LLMs performed similarly to their results in the time-series split method. This result indicates the limitation of smaller parameter models in adapting to different scenarios. Among fine-tuned models, the end-to-end model performed consistently, adapting well to unseen question types and problem-solving styles.
Feedback Generation. We evaluate the quality of these suggestions using five metrics: ROUGE-1, ROUGE-2, ROUGE-L (Lin, 2004), BERTScore (Zhang et al., 2019), and LLM Rating (Liu et al., 2023). The LLM Rating, using GPT-4 as the evaluator, compares the model generated feedback to the ground truth and assigns a score ranging from 0 to 5, with higher scores indicating better quality. The details of the scoring criteria are in Appendix G. The results in Table 6 show that non-fine-tuned language models perform consistently across the time series and leave-one-out splits. Their inability to effectively utilize additional rationales for math problems limits their ability to generate accurate suggestions.
In our fine-tuned models, the single-task and multi-task models performed poorly, rarely identifying errors or generating feedback. Although the end-to-end models achieved ROUGE and BERTScore results comparable to Llama3 70B, they received lower LLM ratings. This represents the complexity of feedback generation, which requires larger models with stronger reasoning and language generation capabilities. Even with fine-tuning, Llama3 8B appears to have reached its performance limit. We also explored the capabilities of o1-mini but conducted only partial experiments due to high financial costs. Main experiments for each task were completed, excluding the leave-one-out and without rationale settings. Results and discussion are in Appendix H.
7 Discussion
7.1 Relationship Between Problem-Solving Ability and Grading Ability
One concern about using LLMs in the math tutoring system is that they may act differently in problem-solving and grading students’ answers. To address this issue, we investigate the correlation between LLMs’ mathematical problem-solving abilities and their accuracy in grading students’ processes. [^13] The model’s performance is presented in Appendix I. A Chi-square test was conducted using predictions from the time-series split setting, excluding rationales to rely solely on the model’s intrinsic capabilities. Since the Chi-square test requires categorical variables, we defined task-specific criteria for statistical validity. Answer accuracy assessment is a binary classification task, allowing direct Chi-square application. Error identification is an extraction task where only exact matches with the ground truth are correct; all others are incorrect, ensuring the results are treated as categorical variables. Feedback generation lacks an absolute correctness measure, so we computed ROUGE-L scores for all generated feedback and set the average score as the threshold, considering feedback correct if it exceeded this value. We report the corresponding $p$-values in Table 7.
Answer Acc. Error Ident. Feedback Gen.
Llama3 8B 1.0000 0.2191 0.8819
Llama3 70B 0.4401 0.3414 0.7407
GPT-3.5 0.8410 0.8538 0.8640
End-to-end 0.3766 0.8852 0.5567
Table 7: Correlation between LLM’s Math Word Problem Solving Ability and Grading Ability.
The results show no significant correlation between the model’s problem-solving ability and its performance in grading tasks. Grading requires more than aligning solutions; it involves mathematical reasoning and step-by-step analysis to assess each part of the student’s process. Effective feedback generation further demands understanding the error, its cause, and providing accurate, context-specific corrections. These findings emphasize the necessity of advanced reasoning and a thorough understanding of mathematical concepts. We also analyzed whether different problem-solving styles among students affect the model’s grading performance. The results show that LLMs perform better when solutions include more textual explanations. Detailed discussions are provided in Appendix J.
7.2 Types of Student Errors that Models Excel at Resolving
To further evaluate the model responses, we selected representative results from Llama3 70B without rationale, GPT-3.5 without rationale, and an end-to-end model with rationale under the time series split setting. Additionally, we included the o1-mini model, which uses few-shot prompts without rationale, to further enhance the comparison and analysis of model performance. The invited three mathematics experts were asked to evaluate the quality of the generated feedback. The evaluation criteria range from 0 (not helpful) to 3 (clear and effective guidance). [^14] The evaluation criteria is elaborated in Appendix K.
All Wrong M. Calc. Incomp. Lack of M. Careless
Llama3 70B 1.61 1.59 1.30 1.33 2.20 2.00
GPT-3.5 1.59 1.46 1.35 1.61 2.30 2.20
o1-mini 1.82 1.79 1.65 1.44 2.55 1.60
End-to-end 1.01 0.81 1.10 0.77 2.00 1.40
Table 8: Human Evaluation of Generated Feedback from Different Models. “Wrong M.,” “Calc.,” “Incomp.,” “Lack of M.,” and “Careless,” correspond to “Wrong Mathematical Operation/Concept,” “Calculation Error,” “Incomplete Answer,” “Lack of Necessary Mathematical Concepts,” and “Careless Error,” respectively.
The results are shown in Table 8.151515The examples and the corresponding analyses are presented in Appendix L. We assessed the inter-rater agreement on the models’ output ratings using Krippendorff’s Alpha, yielding a score of 0.7628, which reflects a moderate level of agreement among the experts. The responses from language models remain below the ideal score of 3. The fine-tuned end-to-end model underperformed compared to larger models. All models excelled in addressing “Lack of Necessary Mathematical Concept” errors. The reason may be that such errors arise when students cannot solve the problem, requiring the model to solve it directly and explain, without analyzing the student’s process.
All models struggled with identifying “Calculation Errors,” often failing even with simple arithmetic mistakes. Interestingly, they performed better on “Wrong Mathematical Operations/Concepts.” Models tended to assume student results were correct and frequently misattributed errors to incorrect methods. In some cases, they even revised correct processes, producing inaccurate feedback. Our fine-tuned end-to-end model performed worse than larger models in “Wrong Mathematical Operations/Concepts,” “Incomplete,” and “Careless Errors,” likely due to its limited reasoning and process analysis capabilities. This led to inaccurate judgments on errors requiring deeper understanding. However, the performance gap narrowed in categories demanding less reasoning, such as “Calculation Errors” and “Lack of Necessary Mathematical Concepts.”
Although GPT-3.5, o1-mini, and Llama3 70B have stronger reasoning and generation capabilities, they struggle to identify student errors and provide focused feedback. Larger models, like o1-mini, often produce verbose responses with inaccuracies, which experts find overwhelming for students and ineffective in addressing specific mistakes. In contrast, our fine-tuned 8B end-to-end model provides concise feedback that directly targets errors but performs poorly on unfamiliar problem types outside its training data. Large models (e.g., GPT-3.5) also tend to solve problems themselves rather than offering adaptive feedback tailored to the student’s reasoning process, even when prompts emphasize error analysis. These results highlight the need for further refinement in LLMs, particularly in handling more nuanced aspects of students’ problem-solving processes.
8 Conclusion
Assessment in mathematics education is critical. However, this grading process is often time-consuming and labor-intensive for teachers, particularly when it requires detailed analysis of each student’s individual problem-solving steps. This study constructed the MathEdu dataset, which includes real student problem-solving processes and teacher feedback for GRE-level math problems. We explore the use of LLMs to automate grading and provide adaptive feedback. While LLMs perform well in identifying correctness and errors, they struggle with generating personalized suggestion. An advanced method to enhance LLMs’ ability to understand and interpret mathematical reasoning is left as our future work.
Limitations
Lack of Student Evaluation. In the current evaluation process, the original students are not involved in reviewing the grading results generated by the LLMs, with only peer students providing feedback. This lack of direct student input may lead to the loss of the most relevant insights. Additionally, since expert annotators evaluating the grading results do not interact with the students, it is difficult to determine to what extent the teacher’s feedback helps students understand their mistakes or improve their learning. In future work, we plan to gather direct opinions from students to determine whether the feedback generated by the model are genuinely helpful to them. By involving students in responding to the model’s grading and recommendations, we can better understand whether the feedback effectively helps students comprehend and correct their mistakes. Such feedback will not only improve the model’s performance but also enhance its applicability in educational contexts.
Limited to Mathematics. While we have conducted several experiments to discuss and analyze the capabilities of LLMs in mathematics tutoring, evaluating their potential advantages and disadvantages in real-world scenarios, our study is constrained to the mathematics domain due to the limitations of our dataset. Research in other fields remains significantly underexplored. In the future, we plan to explore the pedagogical capabilities of LLMs in different educational areas, such as programming instruction, to further uncover their potential in diverse subjects.
Limited Dataset Size. The dataset is constrained to 4,048 entries due to the complexity and effort required for detailed data annotation, limiting its size and potentially reducing its effectiveness in fine-tuning smaller models. Despite its limitations, the dataset offers authentic student problem-solving processes at the GRE difficulty level, accompanied by detailed feedback from expert teachers. It is expected that this resource can support future research in mathematics education, particularly in advancing automated assessment of student answers.
Limited Model and Method Exploration. Our study focused on Llama3 8B, Llama3 70B, and GPT-3.5, with experiments limited to few-shot prompting and LoRA-based fine-tuning. Other models, such as Gemma 2 (Team et al., 2024), Qwen 2.5 (Yang et al., 2024), and Phi-3 (Abdin et al., 2024), as well as newer versions of Llama, remain unexplored. Additionally, our fine-tuning methods did not delve into incorporating students’ individual learning progress, such as tracking their mastered concepts. Despite these limitations, we hope our work can contribute to the growing exploration of LLM applications in mathematics education and encourage further research to refine and expand these methods.
Ethics Statement
To address potential privacy concerns associated with the MathEdu dataset, this section outlines the measures taken to protect annotators’ privacy and the ethical considerations in releasing the dataset. The MathEdu dataset contains problem-solving records annotated by students from diverse academic backgrounds. However, no personally identifiable information, such as names or IDs, was collected. Each record is assigned an anonymized identifier to categorize entries from the same student. While the dataset includes the students’ academic backgrounds (i.e., majors), this information is provided solely for analytical purposes. As highlighted in previous sections, understanding how problem-solving styles vary across mathematical abilities and expertise is a valuable area of research. All student annotators were informed that their personal data would remain confidential and that only their academic background would be disclosed. This consent was obtained prior to participation in the annotation task. The dataset’s structure ensures that individual privacy is preserved while facilitating research in educational applications. In alignment with ethical guidelines, the dataset will only be released for research purposes.
Appendix A Math Word Problem Dataset Selection
Existing open datasets in mathematical domains include datasets such as MAWPS Koncel-Kedziorski et al. (2016), Math23k Wang et al. (2017), SVAMP Patel et al. (2021), and MathQA (Amini et al., 2019). These datasets provide numerous mathematical problems and reference rationales. To gather high-quality student processes, we have two main requirements for the mathematical problems in our dataset:
  1. 1.
    Problem Difficulty: We aim for the dataset to include both problems that can be solved with brief processes and those that are challenging and require complex solution paths. This varying levels of difficulty allow us to obtain diverse problem-solving approaches and explore effective methods for assessing the correctness of these differing strategies.
  2. 2.
    Problem Diversity: We seek a dataset which has multiple domains such as probability, geometry, and algebra. This diversity enables us to comprehensively assess students’ mathematical abilities and further explore techniques applicable to grading problems across different domains.
Considering these factors, we utilize the MathQA dataset (Amini et al., 2019). MathQA includes problems from six distinct domains, covering a range of difficulties suitable for GRE-level problems. Furthermore, MathQA not only provides answers to the problems but also includes rationales, which are correct solution processes that aid teachers in evaluating student process.
Appendix B Supplementary Details for Dataset Construction
B.1 Annotation Guidelines
Table 9 presents the guidelines we used for collecting students’ math word problem solving results. The guidelines begin by explaining the rationale for annotating answers and providing instructions for submitting their work, whether by solving problems on a tablet or solving on paper and uploading a photo. We emphasize the importance of clear annotations and encourage students to include explanations in their problem-solving process. Additionally, we provide examples of proper annotation, and finally, we give instructions on how to handle problems they find unclear.
Student Major Total General Gain Physics Geometry Probability Other
Student 1 Applied Mathematics 683 276 (40.41%) 132 (19.33%) 182 (26.64%) 49 (7.17%) 7 (1.02%) 37 (5.42%)
Student 2 Finance 685 297 (43.36%) 133 (19.42%) 188 (27.45%) 39 (5.69%) 4 (0.58%) 24 (3.50%)
Student 3 Japanese 678 281 (41.45%) 136 (20.06%) 180 (26.55%) 42 (6.19%) 7 (1.03%) 32 (4.72%)
Student 4 Information Management 660 268 (40.60%) 140 (21.21%) 167 (25.30%) 46 (6.97%) 8 (1.21%) 31 (4.70%)
Student 5 Mathematics Education 682 275 (40.32%) 139 (20.38%) 186 (27.27%) 41 (6.01%) 8 (1.17%) 33 (4.84%)
Student 6 Physics 660 262 (39.70%) 127 (19.24%) 175 (26.52%) 55 (8.33%) 7 (1.06%) 34 (5.15%)
Table 10: Student Majors and Distribution of Questions Answered.
Error Type Explanation
Wrong Mathematical Operation/Concept Student applies an incorrect mathematical operation or uses an inappropriate mathematical concept to solve a problem.
Calculation Error Mistakes in calculations, such as errors in solving equations, arithmetic mistakes, and incorrect unit conversions.
Incomplete Answer Student used a correct formula or procedure but did not complete it.
Careless Error Errors caused by students’ carelessness in answering, including number substitution errors and missing digits.
Lack of Necessary Mathematical Concepts Errors in answering caused by a lack of essential mathematical knowledge or techniques.
Table 11: Error Types and Explanations.
B.2 Academic Backgrounds and Question Distribution of Student Annotators
To collect data on math problem-solving, we invited six university students to participate in the annotation process. Table 10 presents the majors of the student annotators and the distribution of questions they answered. We aimed for diversity by including individuals from different academic backgrounds, such as Japanese Studies and Applied Mathematics, to capture a range of math proficiency levels in the problem-solving process. All annotators involved in the task, including students and mathematics experts serving as teachers, were compensated based on the minimum hourly wage standards of their respective countries.
Appendix C Supplementary Information for Each Error Category
Detailed descriptions of these error categories are provided in Table 11. The following are examples of each error category.
Wrong Mathematical Operation/Concept: This category refers to errors where the student applies an incorrect mathematical operation or employs an inappropriate mathematical concept when attempting to solve a problem. Such mistakes often occur when students misunderstand the nature of the mathematical task or misinterpret key elements of the problem. These errors can stem from a variety of issues, including a failure to correctly identify the mathematical procedure required or the application of an irrelevant concept to the problem at hand. Additionally, this category encompasses cases where the student misinterprets critical keywords or phrases, leading to the incorrect selection of operations. For example, misunderstanding terms such as “per,” “rate,” or “difference” can result in choosing the wrong formula or approach. Furthermore, errors in selecting or using relevant information from the problem, either by focusing on irrelevant data or neglecting essential variables, fall under this category. Such errors reflect deeper issues in comprehension or the application of mathematical principles. By recognizing and addressing these types of mistakes, it becomes possible to better understand the student’s reasoning process and to provide more targeted feedback for improving their problem-solving abilities.
In Table 12, we provide an example of this type of error. The problem involves two trains of the same length traveling in the same direction, with the speeds of each train and the time it takes for the faster train to catch up to the slower one provided. The question asks for the length of the trains. While the student correctly calculated that the distance covered by the faster train during the overtaking process was 400 meters, they failed to account for the fact that the faster train needs to cover the combined length of both trains to complete the overtaking maneuver, which led to the incorrect answer.
Problem: Two trains of equal length are running on parallel lines in the same directions at 46 km/hr and 36 km/hr. The faster train passes the slower train in 144 seconds. The length of each train is:
Student Process: $46-36=10$ $10\,\mathrm{km/hr}\times\frac{5}{18}=\frac{100}{36}\,\mathrm{m/s}$ $\frac{100}{36}\times 144=400\,\mathrm{m}$
Error Type: Wrong Mathematical Operation/Concept
Error Equation: $\frac{100}{36}\times 144=400\,\mathrm{m}$
Teacher Feedback: To overtake the other train, you need to travel the combined length of both trains. Since both trains are of the same length, you need to divide by 2 to get the answer.
Table 12: Example of Wrong Mathematical Operation/Concept
Calculation Error: This category captures mistakes made by students during the computational phases of problem-solving, where the execution of arithmetic or algebraic operations is incorrect. Such errors can range from simple miscalculations in basic arithmetic to more complex mistakes involving the manipulation of algebraic expressions or functions. These errors do not typically stem from a misunderstanding of the problem’s structure or concept but rather from faulty execution during the calculation process. Table 13 reports an example of this type of error. The student made a calculation mistake when evaluating $\frac{825-750}{5}$, resulting in an answer of 25 instead of the correct value of 15, leading to an incorrect answer for the problem.
Problem: at what rate percent on simple interest will rs . 750 amount to rs . 825 in 5 years ?
[…]
Error Type: Calculation Error
Error Equation: $\frac{825-750}{5}=25$ Teacher Feedback: Incorrect calculation. The correct answer is 15.
Table 13: Example of Calculation Error
Incomplete Answer: The “Incomplete Answer” category refers to instances where the student begins solving a problem using the correct formula or procedure but does not carry the solution through to completion. Although the initial steps of the problem-solving process may be accurate and aligned with the required methodology, the student ultimately halts the process prematurely. This may result from a partial understanding of the problem or a lack of familiarity with subsequent steps needed to finalize the solution. In Table 14, an example of this error is illustrated. The problem asks for the average number of apples sold per hour over a two-hour period. However, the student calculated only the total number of apples sold during those two hours and stopped there, failing to continue to calculate the average, which led to an incorrect answer.
Problem: maria sold 10 kg of apples in her first hour at the market , but only 2 kg of apples in the second hour . on average , how many kg of apples did she sell in two hours at the market ?
Student Process: 10+2=12
Error Type: Incomplete Answer
Error Equation: None
Teacher Feedback: After calculating the total number of apples sold in two hours, you still need to divide by the time to get the average sales per hour. Therefore, the answer is 12/2 = 6.
Table 14: Example of Incomplete Answer
Careless Error: The “Careless Error” category encompasses mistakes that arise not from a misunderstanding of the mathematical concepts or procedures, but rather from inattentiveness or lapses in concentration during the problem-solving process. These errors often result from avoidable mistakes, such as substituting incorrect numbers into a formula or writing the wrong number. Table 15, shows an example of this type of error. The problem provides the number 78; however, the student mistakenly interprets this number as 70 during the calculations, leading to an incorrect result.
Problem: Peter’s average (arithmetic mean) test score on 4 tests is 78. What must be Peter’s score on the 5th test for his average score on the 5 tests to be 80?
Student Process: $80\times 5-70\times 4=120$
Error Type: Careless Error
Error Equation: $80\times 5-70\times 4=120$
Teacher Feedback: Carelessly misreading the numbers in the problem, the original average was 78 instead of 70. Therefore, the correct process is 805 - 784 = 88.
Table 15: Example of Careless Error
Lack of Necessary Mathematical Concepts: The “Lack of Necessary Mathematical Concepts” category refers to errors that occur when students lack fundamental mathematical knowledge or techniques needed to solve a problem. These errors often result in students being unable to attempt the problem at all, as they may not possess the requisite understanding of essential concepts such as fractions, percentages, or algebraic expressions. Table 16 presents an example of this type of error. The question involved calculating compound interest, but the student responded with “Do not know how to calculate compound interest.” indicating a lack of knowledge on how to perform this calculation. As a result, the student was unable to arrive at the correct solution for the problem.
Problem:If \$5000 is invested in an account that earns 12% interest compounded semi-annually, then the interest earned after one year would be how much greater than if the \$5000 had been invested at 8% simple yearly interest?
Student Process: Do not know how to calculate compound interest.
Error Type: Lack of Necessary Mathematical Concepts
Error Equation: None
Teacher Feedback: When calculating compound interest, you also need to consider the interest generated from the previous year’s interest. The formula is $A=P\left(1+\frac{r}{n}\right)^{nt}$ where $P$ is the principal, $r$ is the annual interest rate, $n$ is the number of times interest is compounded per year, and $t$ is the number of years the money is invested or borrowed. Therefore, the calculation for compound interest in this question is $A=5000\left(1+0.12\right)^{2}.$
Table 16: Example of Lack of Necessary Mathematical Concepts
Appendix D Details of Fine-Tuning Methods
We employ LoRA fine-tuning to train an LLM $\mathcal{M}$, parameterized by $\Phi$. The training data used in LoRA fine-tuning is denoted as $\mathcal{Z}=\{(x_{i},y_{i})\}_{i=1,\ldots,N}$. Given task-specific parameter increment $\Delta\Phi=\Delta\Phi(\Theta)$, we optimize over $\Theta$:
$\max_{\Theta}\sum_{(x,y)\in\mathcal{Z}}\sum_{t=1}^{|y|}(\log p_{{\Phi}_{0}+\Delta\Phi(\Theta)}(y_{t}|x,y_{<t})).$ (1)
Single-task and multi-task with rationale: In these two approaches, the training data used in LoRA fine-tuning is denoted as $\mathcal{Z}=\{(x_{i},c_{i}),(x_{i},E_{i}),(x_{i},T_{i})\}_{i=1,\ldots,N},$ where $x_{i}=\mathcal{P}_{A}(q_{i},p_{i},r_{i})$. Here, $\mathcal{P}_{A}$ represents the prompt that takes $q_{i}$, $r_{i}$, and $p_{i}$ to generate grading results.
Single-task and multi-task without rationale: Similar to the “Single-task and multi-task with rationale” approaches, these methods also use the same training approach but exclude the rationale in the input data. The training data used in LoRA fine-tuning is denoted as $\mathcal{Z}=\{(x_{i},c_{i}),(x_{i},E_{i}),(x_{i},T_{i})\}_{i=1,\ldots,N},$ where $x_{i}=\mathcal{P}_{B}(q_{i},p_{i})$. Here, $\mathcal{P}_{B}$ represents the prompt that takes $q_{i}$ and $p_{i}$ to generate grading results.
End-to-end with rationale: In this approach, the training data used in LoRA fine-tuning is denoted as $\mathcal{Z}=\{(x_{i},y_{i})\}_{i=1,\ldots,N},$ where $x_{i}=\mathcal{P}_{C}(q_{i},p_{i},r_{i})$ and $y_{i}=(c_{i},E_{i},T_{i})$. Here, $\mathcal{P}_{C}$ represents the prompt that takes $q_{i}$, $r_{i}$, and $p_{i}$ to generate grading results.
End-to-end without rationale: Similar to the “End-to-end with rationale” approach, this method also uses an end-to-end training approach but excludes the rationale in the input data. The training data used in LoRA fine-tuning is denoted as $\mathcal{Z}=\{(x_{i},y_{i})\}_{i=1,\ldots,N},$ where $x_{i}=\mathcal{P}_{D}(q_{i},p_{i})$ and $y_{i}=(c_{i},E_{i},T_{i})$. Here, $\mathcal{P}_{D}$ represents the prompt that takes $q_{i}$ and $p_{i}$ to generate grading results.
Appendix E Input Format
This section presents the few-shot prompts and fine-tuning prompts used for the answer accuracy assessment task. The prompt for rating the generated feedback is also included. Table 17 presents the few-shot prompts. Since we examined the models both with and without rationale, the prompts are presented in two formats. In these prompts, “few-shot examples” denotes the position for examples, where six samples are randomly selected from the training dataset, comprising three correct and three incorrect responses.
Without Rationale With Rationale
You are a math teacher. According to the [Question], please indicate whether the [Student’s Answer] is correct or not. If the [Student’s Answer] is incorrect, identify where the student went wrong and provide explanations and advice. For correct answers, respond: The student’s answer is correct. For incorrect answers, respond: The student’s answer is incorrect. [wrong equation] ... [Teacher’s explanations and advice] ... {few-shot examples} [Question] : {question} Option: {options} [Student’s Answer] : {student process} You are a math teacher. According to the [Question] and [Rationale], please indicate whether the [Student’s Answer] is correct or not. If the [Student’s Answer] is incorrect, identify where the student went wrong and provide explanations and advice. For correct answers, respond: The student’s answer is correct. For incorrect answers, respond: The student’s answer is incorrect. [wrong equation] ... [Teacher’s explanations and advice] ... {few-shot examples} [Question] : {question} Option: {options} [Rationale] : {rationale} [Student’s Answer] : {student process}
Table 17: Few-Shot Prompts for Grading Student Problem-Solving Results.
Table 18 presents the end-to-end prompts. Since the fine-tuned models are trained and inferred using a zero-shot approach, there is no “few-shot examples” field included. The prompt content has been adjusted to meet the format required by Llama. Table 19 presents the prompts used for the single-task and multi-task models. Because we separated the tasks of answer accuracy assessment, problem-solving Error Identification, and feedback generation, we created three distinct prompts for each task.
Without Rationale With Rationale
### Instruction: You are a math teacher. According to the [Question], please indicate whether the [Student’s Answer] is correct or not. If the [Student’s Answer] is incorrect, identify where the student went wrong and provide explanations and advice. For correct answers, respond: The student’s answer is correct. For incorrect answers, respond: The student’s answer is incorrect. [wrong equation] ... [Teacher’s explanations and advice] ... ### Input: [Question] : {question} Option: {options} [Student’s Answer] : {student process} ### Response: ### Instruction: You are a math teacher. According to the [Question] and [Rationale], please indicate whether the [Student’s Answer] is correct or not. If the [Student’s Answer] is incorrect, identify where the student went wrong and provide explanations and advice. For correct answers, respond: The student’s answer is correct. For incorrect answers, respond: The student’s answer is incorrect. [wrong equation] ... [Teacher’s explanations and advice] ... ### Input: [Question] : {question} Option: {options} [Rationale] : {rationale} [Student’s Answer] : {student process} ### Response:
Table 18: End-to-End Fine-Tuning Prompt for Grading Student Problem-Solving Results.
Category Without Rationale With Rationale
Answer Accuracy Assessment ### Instruction: You are a math teacher. According to the [Question], please indicate whether the [Student’s Answer] is correct or not. For correct answers, respond: The student’s answer is correct. For incorrect answers, respond: The student’s answer is incorrect. ### Input: [Question] : {question} Option: {options} [Student’s Answer] : {student process} ### Response: ### Instruction: You are a math teacher. According to the [Question] and [Rationale], please indicate whether the [Student’s Answer] is correct or not. For correct answers, respond: The student’s answer is correct. For incorrect answers, respond: The student’s answer is incorrect. ### Input: [Question] : {question} Option: {options} [Rationale] : {rationale} [Student’s Answer] : {student process} ### Response:
Problem-Solving Error Identification ### Instruction: You are a math teacher. According to the [Question], please indicate whether the [Student’s Answer] contains any wrong equations. For correct answers, respond: The student’s answer doesn’t contain wrong equations. For incorrect answers with wrong equations, respond: [wrong equation] ... ### Input: [Question] : {question} Option: {options} [Student’s Answer] : {student process} ### Response: ### Instruction: You are a math teacher. According to the [Question] and [Rationale], please indicate whether the [Student’s Answer] contains any wrong equations. For correct answers, respond: The student’s answer doesn’t contain wrong equations. For incorrect answers with wrong equations, respond: [wrong equation] ... ### Input: [Question] : {question} Option: {options} [Rationale] : {rationale} [Student’s Answer] : {student process} ### Response:
Feedback    Generation ### Instruction: You are a math teacher. According to the [Question], please indicate whether the [Student’s Answer] is correct or not. If the [Student’s Answer] is incorrect, explain why the student went wrong and provide advice on how to correct their mistake. For correct answers, respond: The student’s answer is correct and doesn’t need advice. For incorrect answers, respond: [Teacher’s explanations and advice] ... ### Input: [Question] : {question} Option: {options} [Student’s Answer] : {student process} ### Response: ### Instruction: You are a math teacher. According to the [Question] and [Rationale], please indicate whether the [Student’s Answer] is correct or not. If the [Student’s Answer] is incorrect, explain why the student went wrong and provide advice on how to correct their mistake. For correct answers, respond: The student’s answer is correct and doesn’t need advice. For incorrect answers, respond: [Teacher’s explanations and advice] ... ### Input: [Question] : {question} Option: {options} [Rationale] : {rationale} [Student’s Answer] : {student process} ### Response:
Table 19: Single-Task and Multi-Task Learning Prompts for Grading Student Problem-Solving Results.
Table 20 presents the prompts used for rating the generated feedback, where we first outline the rating criteria, with scores ranging from 0 to 5, indicating poor to excellent. Additionally, we specify that the output should follow the format: “The rating for this correct suggestion is:” to facilitate the result collection process.
Given a problem, student’s process and the teacher’s manual suggestions, please rate the following feedback generated by different models on a scale from 0 to 5. The criteria for the rating are:
0: Completely irrelevant or nonsensical suggestions. 1: Very poor suggestions with little relevance or helpfulness. 2: Poor suggestions with some relevance but mostly unhelpful. 3: Average suggestions that are somewhat relevant and somewhat helpful. 4: Good suggestions that are relevant and helpful but could be improved. 5: Excellent suggestions that are highly relevant, helpful, and well-aligned with the teacher’s manual suggestions.
Please only answer in the following format:
The rating for this feedback is:
Table 20: Rating Prompt.
Train Validation Test
Time Series Split 2,836 609 603
Leave-One-Out S1 3,028 337 683
S2 3,026 337 685
S3 3,033 337 678
S4 3,049 339 660
S5 3,029 337 682
S6 3,049 339 660
Table 21: Data Count for Various Data Splitting Strategies.
Appendix F Data Splitting Strategies and Statistics
Table 21 presents the data counts for the time-series split and leave-one-out methods for different students in the train, validation and test sets.
Appendix G Distance Metric
The exact match metric is overly strict, so we also consider cases where the model’s predicted errors slightly deviate from the annotated errors. For example, minor character differences, though not entirely correct, may still be acceptable. Similarly, the model might predict the step immediately before or after the actual error. The distance metric is designed to distinguish these minor deviations from predictions that are entirely misaligned with the annotated errors, providing a more comprehensive evaluation of the model’s ability to detect errors accurately.
To evaluate the discrepancy between the predicted erroneous equations $P_{i}$ and the annotated erroneous equations $G_{i}$, we compute the distance $D(P_{i},G_{i})$ for each pair. The positions of these equations within the student’s solution are represented by $P_{\text{start}}$ and $P_{\text{end}}$ for $P_{i}$, and $G_{\text{start}}$ and $G_{\text{end}}$ for $G_{i}$. Specifically, $P_{\text{start}}$ is the starting character position of $P_{i}$, and $P_{\text{end}}$ is the ending position. Similarly, $G_{\text{start}}$ is the starting position of $G_{i}$, and $G_{\text{end}}$ is the ending position.
A distance of 0 is assigned when the model accurately identifies the error. If the model incorrectly classifies an incorrect answer as correct, a penalty of 127 is applied, representing the average length of the student’s solution. Specifically, the calculation of $D(P_{i},G_{i})$ is defined as follows:
  1. 1.
    If both $P_{i}$ and $G_{i}$ are “None” (indicating no errors in the equation), the distance $D(P_{i},G_{i})=0$.
  2. 2.
    If only one of $P_{i}$ or $G_{i}$ is “None,” the distance is set to the penalty value $D_{\text{penalty}}$.
  3. 3.
    If the start and end positions of $P_{i}$ and $G_{i}$ perfectly match, the distance $D(P_{i},G_{i})=0$.
  4. 4.
    When $P_{i}$ and $G_{i}$ do not overlap, the distance is the character difference between their boundaries:
    • If $P_{\text{end}}\leq G_{\text{start}}$, then $D(P_{i},G_{i})=G_{\text{start}}-P_{\text{end}}$.
    • If $G_{\text{end}}\leq P_{\text{start}}$, then $D(P_{i},G_{i})=P_{\text{start}}-G_{\text{end}}$.
  5. 5.
    When $P_{i}$ and $G_{i}$ partially overlap, the distance is the sum of the non-overlapping character differences:
    • If $G_{\text{start}}\leq P_{\text{start}}\leq P_{\text{end}}\leq G_{\text{end}}$, then $D(P_{i},G_{i})=(P_{\text{start}}-G_{\text{start}})+(G_{\text{end}}-P_{\text{end}})$.
    • If $P_{\text{start}}\leq G_{\text{start}}\leq G_{\text{end}}\leq P_{\text{end}}$, then $D(P_{i},G_{i})=(G_{\text{start}}-P_{\text{start}})+(P_{\text{end}}-G_{\text{end}})$.
The distance is calculated for each pair of $P_{i}$ and $G_{i}$, and the minimum distance is taken as the distance for the corresponding equation pair. Formally, the distance is measured as follow:
$D(P_{i},G_{i})=$ (2)
$\begin{cases}0,&\text{if both }P_{i}\text{ and }G_{i}\text{ are ``None''}\\ \text{min}(D_{\text{penalty}},\text{ret\_dis}),&\text{if one of }P_{i}\text{ or }G_{i}\text{ is ``None''}\\ 0,&\text{if }P_{i}\text{ and }G_{i}\text{ perfectly match}\\ G_{\text{start}}-P_{\text{end}},&\text{if }P_{\text{end}}\leq G_{\text{start}}\\ P_{\text{start}}-G_{\text{end}},&\text{if }G_{\text{end}}\leq P_{\text{start}}\\ (P_{\text{start}}-G_{\text{start}})+(G_{\text{end}}-P_{\text{end}}),&\text{if }G_{\text{start}}\leq P_{\text{start}}\leq P_{\text{end}}\leq G_{\text{end}}\\ (G_{\text{start}}-P_{\text{start}})+(P_{\text{end}}-G_{\text{end}}),&\text{if }P_{\text{start}}\leq G_{\text{start}}\leq G_{\text{end}}\leq P_{\text{end}}\\ \text{min}(D_{\text{penalty}},\text{ret\_dis}),&\text{otherwise.}\end{cases}$
$\text{Distance}=\frac{1}{n}\sum_{i=1}^{n}D(P_{i},G_{i})$ (3)
where $n$ is the number of instance.
Appendix H o1-mini Result
The o1-mini model, developed by OpenAI, is renowned for its strong mathematical reasoning capabilities. To evaluate its performance, we conducted experiments using a time series split setting, where inference was performed using few-shot prompts without rationales. We analyzed its performance across three tasks and compared the results with our fine-tuned end-to-end model.
The results of the answer accuracy assessment task are presented in Table 22. Notably, o1-mini demonstrates exceptional mathematical capabilities, achieving an overall accuracy of 92.54% in identifying the correctness of the student’s answers. This surpasses the best-performing end-to-end model with rationale.
The results of the error identification task are shown in Table 23. o1-mini demonstrated stronger mathematical reasoning and understanding abilities, achieving a 41.98% exact match ratio, surpassing the end-to-end model. However, it performed worse in terms of the distance metric. Upon reviewing o1-mini’s outputs, we found that in cases where errors were due to unfinished problems, the student’s calculations were correct. Teachers typically did not mark any formulas as incorrect. Despite this, o1-mini still flagged incorrect formulas in these cases, which introduced significant penalties in the distance metric, leading to a poorer performance in this aspect.
The results of the feedback generation task are shown in Table 24. R-1, R-2, R-L, and Rating represent ROUGE-1, ROUGE-2, ROUGE-L, and LLM Rating metrics, respectively. o1-mini outperformed the end-to-end model in terms of BERTScore. However, its responses were often overly lengthy, generating additional irrelevant information, which impacted the evaluation results, and limited it performance compared to GPT-3.5. In the LLM rating conducted by using GPT-4, o1-mini received an almost perfect score of 4.70. By contrast, in the evaluation conducted by human annotators (i.e., Table 8), o1-mini’s performance was less satisfactory in generating adaptive feedback, as its responses were often overly lengthy and lacked focus, reducing their effectiveness in addressing specific student errors.
All General Gain Physics Geo. Prob. Other
o1-mini w/o $r$ 92.54% 92.24% 92.79% 92.40% 90.00% 93.75% 96.97%
End-to-end w/o $r$ 87.73% 84.05% 93.69% 90.06% 90.00% 62.50% 90.91%
End-to-end w/ $r$ 91.54% 88.79% 95.50% 93.57% 95.00% 62.50% 96.97%
Table 22: o1-mini’s Results of Answer Accuracy Assessment.
EM ($\uparrow$) Distance ($\downarrow$)
o1-mini w/o $r$ 41.98% 56.30
End-to-end w/o $r$ 30.25% 71.48
End-to-end w/ $r$ 40.12% 51.78
Table 23: o1-mini’s Results of Problem-Solving Error Identification.
R-1 R-2 R-L BERTScore Rating
o1-mini w/o $r$ 0.1956 0.0633 0.1839 0.7986 4.7
End-to-end w/o $r$ 0.1862 0.0751 0.1761 0.5327 0.99
End-to-end w/ $r$ 0.2842 0.1418 0.2703 0.6786 1.91
Table 24: o1-mini’s Results of feedback generation.
Appendix I Mathematical Problem-Solving Ability of the Model
To evaluate the problem-solving capabilities of the models used in the baseline with respect to the questions in our dataset, we employed the test set from the time series split method. The results, presented in Table 25, show that even the smallest parameter model, Llama 8B, achieved an overall accuracy of 75.95%. In contrast, Llama 80B attained an impressive overall accuracy of 92.37%, surpassing all of our students. This demonstrates that the LLMs we utilized are capable of understanding and solving the problems present in this task.
Category Questions LLaMA3 8B LLaMA3 70B GPT-3.5
All 603 75.95% 92.37% 82.59%
General 232 75.43% 90.51% 84.91%
Gain 111 80.18% 93.69% 88.29%
Physics 171 73.68% 95.32% 78.36%
Geometry 40 77.50% 87.50% 77.50%
Probability 16 75.00% 100.00% 81.25%
Other 33 75.76% 87.88% 75.76%
Table 25: Performance of Different Models Across Various Problem Categories.
S1 S2 S3 S4 S5 S6
Llama3 8B w/o $r$ 59.59% 70.36% 61.50% 67.88% 66.72% 55.15%
w/ $r$ 62.37% 72.26% 64.90% 71.06% 73.90% 66.97%
Llama3 70B w/o $r$ 84.77% 88.91% 84.96% 82.88% 87.24% 82.73%
w/ $r$ 87.70% 86.13% 83.78% 81.67% 81.67% 83.79%
GPT-3.5 w/o $r$ 40.85% 47.45% 45.28% 36.36% 51.32% 28.94%
w/ $r$ 50.07% 65.99% 50.74% 47.12% 52.05% 29.55%
Single-task w/o $r$ 20.06% 87.88% 85.25% 67.73% 85.19% 85.45%
w/ $r$ 93.85% 94.01% 92.04% 89.85% 92.08% 92.88%
Multi-task w/o $r$ 0.00% 30.80% 66.08% 70.61% 90.91% 0.00%
w/ $r$ 89.75% 88.03% 47.20% 0.00% 85.34% 18.33%
End-to-end w/o $r$ 87.12% 57.96% 85.55% 87.73% 90.03% 89.09%
w/ $r$ 94.58% 95.33% 90.86% 89.85% 96.48% 91.67%
Table 26: Model Performance on Answer Accuracy Assessment Across Different Students.
Appendix J Impact of Variations in Student Problem-Solving Styles on Model Grading
The problem-solving results in our dataset exhibit considerable diversity among students, as detailed in Table 3. Students 2, 3, and 4 provided more detailed and complex solutions, often including natural language explanations. In contrast, Students 1, 5, and 6 used fewer equations and offered more abbreviated responses, with Students 1 and 6 almost entirely omitting natural language descriptions of their processes. The variation in student response styles raises the question of whether this diversity impacts the models’ grading performance.
We can examine this question by looking at the experimental results in Table 26 under the leave-one-out data split method. For the non-finetuned models, the number of parameters plays a significant role in their ability to interpret and assess students’ problem-solving results. The Llama3 8B model, with fewer parameters, struggled to accurately grade the concise and less descriptive responses provided by Students 1 and 6, resulting in a lower performance. Conversely, the Llama3 70B model, although still affected by these abbreviated responses, demonstrated greater resilience due to its enhanced comprehension capabilities. Similarly, GPT-3.5, despite lower overall accuracy due to its more stringent evaluation of student processes, exhibited comparable trends across the different students, aligning with Llama3 in terms of accuracy variation.
For the fine-tuned models, those trained with the multi-task learning method demonstrated less robustness. We found that these models occasionally failed to provide any output when confronted with particularly brief responses, leading to instances of zero accuracy. Moreover, the fine-tuned models showed reduced performance when assessing more detailed and complex solutions, such as those from Student 2, likely due to limitations in their ability to process and understand intricate reasoning. However, when models were allowed to reference rationales during answer accuracy assessment, their overall performance improved, and discrepancies in grading accuracy across different problem-solving styles were mitigated. That means the incorporation of rationales can assist models in navigating the challenges posed by varied student responses, enabling them to more accurately determine whether a student’s problem-solving process is correct.
To sum up, the experiment indicates that the diversity in students’ problem-solving styles affect the models’ answer accuracy assessment performance. If the student’s solution includes more detailed explanations or even natural language descriptions, the LLMs are better able to evaluate the solution accurately, leading to more reliable grading outcomes.
Appendix K Evaluation Criteria for Generated Feedback Quality
The evaluation criteria are based on a 0 to 3 scale, as follows:
  • 0: The model’s feedback did not help the student at all.
  • 1: The model’s feedback pointed in the right direction but contained errors, making it unhelpful to the student.
  • 2: The model’s feedback correctly indicated how to correct the student’s mistake but lacked clarity, potentially causing the student difficulty in understanding.
  • 3: The model’s feedback effectively guided the student on how to correct the mistake, with clear and detailed explanations.
Appendix L Analysis of Model-Generated Feedback
In this section, we present examples of feedback generated by the models to analyze their distinct behaviors. Table 27 is a relatively simple math word problem, specifically asking how many apples Maria sells on average per hour over a two-hour period. The student only calculated the total number of apples sold during the two hours without determining the average, resulting in an incorrect answer.
Llama3 70B accurately pointed out that the student only calculated the total number of apples sold rather than the average and provided guidance on how to proceed with the calculation, offering excellent advice. GPT-3.5 correctly noted that the student calculated the total number of apples sold instead of the average and explained how to perform the subsequent calculations. However, while their suggestions were detailed, the content was overwhelming for the student, making it difficult for them to quickly understand how to correct their answer. On the other hand, our fine-tuned end-to-end model effectively highlighted that the student needed to divide the total by 2 to arrive at the required average number of apples sold, providing valuable suggestion.
In Table 28, another example is presented where the question asks which orientation of a water tank placed in a crate measuring 8 by 12 by 14 feet would yield the largest volume. The tank is to be placed upright, and the student mistakenly used the smallest dimension divided by 2 as the radius, leading to an error. The problem requires consideration of different orientations of the crate, ultimately determining that placing the crate on the 12 by 14 face yields the maximum tank volume, with the correct radius being 6 feet.
Llama3 70B recognized that the student made an error but ended up performing the same calculation as the student, arriving at the same incorrect answer, which was unhelpful. Similarly, GPT-3.5 also failed to consider the varying tank volumes based on different crate orientations. It acknowledged the student’s mistake but provided the same incorrect answer, rendering it similarly unhelpful.
In contrast, o1-mini demonstrated the strongest mathematical reasoning abilities, correctly identifying the solution to the problem. It provided a detailed explanation of each step and clarified the reasons for the student’s mistake. Although the explanation was somewhat lengthy, it still constituted a helpful suggestion. The end-to-end model not only failed to offer a correct correction but also misinterpreted the student’s process, mistakenly assuming the student was calculating the crate’s dimensions. This model performed the worst in this instance.
Problem:maria sold 10 kg of apples in her first hour at the market , but only 2 kg of apples in the second hour . on average , how many kg of apples did she sell in two hours at the market ?
Student Process: 10+2=12
Human: After calculating the total number of apples sold in two hours, you still need to divide by the time to get the average sales per hour. Therefore, the answer is 12/2 = 6.
Llama3 70B: The student has calculated the total amount of apples sold, but not the average. To find the average, divide the total amount by the number of hours: $\frac{(10+2)}{2}=\frac{12}{2}=6.$ GPT-3.5: To find the average, you need to add the total amount of apples sold in both hours and then divide by the total number of hours. In this case, Maria sold 10 kg in the first hour and 2 kg in the second hour, so the total amount of apples sold in two hours is 10 kg + 2 kg = 12 kg. To find the average, you need to divide this total by 2 (the number of hours), so the average amount of apples sold in two hours is 12 kg / 2 = 6 kg. Therefore, Maria sold an average of 6 kg of apples in two hours at the market.
[…]
End-to-end: The question asks for the average number of kilograms of apples sold in two hours, not the total number of kilograms of apples sold in two hours. Therefore, you need to divide the total number of kilograms of apples sold in two hours by 2 to get the average number of kilograms of apples sold in two hours.
Table 27: Comparison of Model-Generated Suggestions on Arithmetic Problem.
Problem:a certain rectangular crate measures 8 feet by 12 feet by 14 feet . a cylindrical gas tank is to be made for shipment in the crate and will stand upright when the crate is placed on one of its six faces . what should the radius of the tank be if it is to be of the largest possible volume ?
Student Process: $\frac{8}{2}=4\;\text{feet}$
Human: The suggestion from the teacher is that simply taking the shortest edge of the crate to calculate the radius is incorrect. It is necessary to consider the different orientations of the crate. Since the gas tank will be placed inside the crate, the radius of the gas tank will depend on the shorter edge of the face on which the crate is placed. By calculating the scenarios for the three different faces, we can determine that placing the gas tank on the 12 feet by 14 feet face will yield the largest volume, calculated as $\left(\frac{12}{2}\right)\times\left(\frac{12}{2}\right)\times\pi\times 8$ Therefore, the radius is 6.
Llama3 70B: The student seems to be trying to find the radius of the cylinder, but the calculation is incorrect. Since the crate measures 8 feet by 12 feet by 14 feet, the largest possible radius of the cylindrical tank would be the smallest dimension of the crate, which is 8 feet, divided by 2, which is 4 feet.
GPT-3.5: To find the largest possible volume for the cylindrical gas tank that can fit inside the rectangular crate, the radius of the tank should be half of the smaller dimension of the crate. In this case, the radius should be half of 8 feet, which is 4 feet.
o1-mini: The student only considered one possible orientation of the crate by taking the dimension 8 feet and dividing it by 2 to get a radius of 4 feet. However, to achieve the largest possible volume for the cylindrical gas tank, it’s essential to evaluate all possible orientations of the crate. Here’s how to approach the problem correctly: 1. Identify all orientations: (a) First Orientation: Placing the crate on the 8 ft $\times$ 12 ft face. Height of cylinder: 14 ft Maximum diameter: 8 ft (since the cylinder must fit within the 8 ft width) Radius: $\frac{8}{2}=4$ ft Volume: $\pi\times 4^{2}\times 14=224\pi$ cubic feet (b) Second Orientation: Placing the crate on the 12 ft $\times$ 14 ft face. Height of cylinder: 8 ft Maximum diameter: 12 ft Radius: $\frac{12}{2}=6$ ft Volume: $\pi\times 6^{2}\times 8=288\pi$ cubic feet (c) Third Orientation: Placing the crate on the 8 ft $\times$ 14 ft face. Height of cylinder: 12 ft Maximum diameter: 8 ft Radius: $\frac{8}{2}=4$ ft Volume: $\pi\times 4^{2}\times 12=192\pi$ cubic feet 2. Determine the optimal orientation: The second orientation (placing the crate on the 12 ft $\times$ 14 ft face) provides the largest volume for the cylinder with a radius of 6 feet. Conclusion:
To maximize the volume of the cylindrical gas tank, the crate should be placed on the 12 ft $\times$ 14 ft face, allowing the tank to have a radius of 6 feet.
End-to-end: The question is misunderstood. The question asks for the radius of the tank, not the length of the crate.
Table 28: Comparison of Model-Generated Suggestions to a Geometric Problem Involving Volume Calculation.
Appendix M Information About Use Of Ai Assistants
We utilized generative AI tools solely for grammar and language refinement. Following the use of this tool, the author(s) carefully reviewed and revised the content as necessary, assuming full responsibility for the final publication.